# Local gauge invariant operator on isometry breaking background

### v1

@@ 1 Introduction
Local field operators are the main building blocks of local quantum field theory, playing a key role in combining quantum mechanics with special relativity successfully (see, e.g., [2, 3] and references therein). Nevertheless, not all of them are physically meaningful when gravity is taken into account : for spacetime to be consistently treated as a dynamical object, physical operator must be invariant under the diffeomorphism. Indeed, it was argued that such gauge invariant operator is typically non-local, which is realized by ‘dressing’ the local operator with the Wilson line connecting the point where the local operator is defined and the boundary [4, 5, 6, 7, 8, 9]. This ‘dressing theorem’ can be shown by observing that if the background has a boundary or an asymptotic region, the diffeomorphism generator $\Phi$ contains a boundary term $\Phi_{\partial}$ of ${\cal O}(\kappa^{-1})$ where $\kappa=(8\pi G)^{1/2}$, thus can be written as $\Phi=\Phi_{\rm bulk}+\Phi_{\partial}$. The local operator $O$ in the bulk commutes with the boundary term ($O,\Phi_{\partial}]=0$) while the commutator $O,\Phi_{\rm bulk}]$ of ${\cal O}(\kappa^{0})$ is typically nonzero. Then $O$ can be promoted to the gauge invariant operator by corrected by the Wilson line $\kappa O^{\prime}$ connecting the bulk and the boundary, as the commutator $\kappa O^{\prime},\Phi_{\partial}]$ of ${\cal O}(\kappa^{0})$ is nonvanishing but identical to $O,\Phi_{\rm bulk}]$. In the end, the dressed operator $O+\kappa O^{\prime}$ is gauge invariant and nonlocal.
@@
However, if all the gauge invariant operators are nonlocal, it is challenging to explain several quantum gravity phenomena. First, it was recently suggested that the black hole information paradox can be resolved by appearance of the local region called the ‘island’ in the black hole interior after the Page time [10, 11, 12, 13, 14] (see also [15, 16] for reviews). As the operators localized on the island can be reconstructed from the operators for the Hawking radiation far away from the black hole, the entropy of the Hawking radiation does not increase indefinitely, which is consistent with unitarity. But in the absence of local gauge invariant operator, the operators localized on the island are difficult to be well defined. Second, whereas de Sitter (dS) space well describes the observed accelerating expansion of our universe [17, 18], it does not have a boundary, so the Wilson line connecting the bulk and the boundary is not defined.
@@
Regarding the first issue of the island, it has been suggested to construct the local operator by breaking the diffeomorphism invariance explicitly, resulting in a massive graviton [19, 20, 21, 22, 23]. This is motivated by the fact that when we study the island of the anti-de Sitter (AdS) black hole in the context of AdS/CFT correspondence [24, 25, 26], the boundary of AdS space is assumed to be attached to the ‘bath’ in which gravity is not dynamical and the old Hawking radiation is contained. Whereas it is a convenient way to separate the Hawking radiation from dynamics of gravity, since energy is allowed to pass through the border between AdS space and the bath, the gravitational energy-momentum tensor is not conserved. In other words, the diffeomorphism invariance is broken by the boundary condition. This can be modelled by introducing the compensating vector field $V_{\mu}$ which transforms as $V_{\mu}\to V_{\mu}-\epsilon_{\mu}$ under the diffeomorphism $x^{\mu}\to x^{\mu}+\epsilon^{\mu}$ before being fixed to some specific function of $x^{\mu}$. Then $V_{\mu}$ behaves like the Goldstone vector field : it combines with the fluctuation of the metric $h_{\mu\nu}$ to form the would-be gauge invariant combination $\tilde{h}_{\mu\nu}=h_{\mu\nu}+\nabla_{\mu}V_{\nu}+\nabla_{\nu}V_{\mu}$, allowing the graviton mass term $(M^{4}/4)(\tilde{h}_{\mu\nu}\tilde{h}^{\mu\nu}-\tilde{h}^{\mu}_{\mu}\tilde{h}^{\nu}_{\nu})$ in the Lagrangian density [23]. This corresponds to the imitation of (the tensor version of) the Stückelberg (or the Higgs in a broad sense) mechanism, in the vector version of which gauge invariance is spontaneously broken as the gauge boson becomes massive by absorbing a scalar transforming nonlinearly under the gauge transformation. Moreover, any local operator $O(x)$ can be promoted to the local would-be gauge invariant operator O(xV(x))O(x-V(x)) which is expected to be the form of operators localized on the island [27].
@@
Meanwhile, it was pointed out in [28] that the tension between the gauge invariance and the locality discussed so far is present only if the background has a symmetry called isometry. Intuitively, we may understand that the nonlocal connection between the bulk and the boundary is closely related to the isometry from the fact that any fluctuation breaking the isometry raises the energy, which must be reflected in the fluctuation of the asymptotic charge like the Arnowitt-Deser-Misner (ADM) Hamiltonian. In contrast, in the absence of the isometry, the asymptotic charge may not be affected by the fluctuation in the bulk. Technically, the nonexistence of the local operator on the background having the isometry can be explained by the absence of the clock and the rod. To see this more concretely, suppose the background has a timelike isometry. In this case, one finds a time direction along which the metric does not change, hence the classical solution for the matter field is independent of the corresponding time coordinate $t$ as well. Then there is no way to distinguish some particular time slice from another, which is said to be the absence of the clock. In the same way, the spacelike isometry prevents distinguishing a particular spatial slice from another, then we say that the rod does not exist.
@@
In fact, the idea that the local gauge invariant operator can be constructed by combining the matter and the metric fluctuations goes back to [29]. It was immediately used to describe the cosmological fluctuations around the Friedmann-Robertson-Walker (FRW) background which may not have a boundary, resolving the second issue [30, 31, 32, 33, 34, 35]. In particular, the curvature perturbation, the local gauge invariant operator made up of the scalar fluctuation and the trace of the spatial metric fluctuation, well describes the primordial inhomogeneity generated in the inflationary era, in which the background geometry is given by quasi-dS space where the dS isometry is slightly broken [36, 37]. The Stückelberg mechanism for the curvature perturbation must be distinguished from the imitated Stückelberg mechanism with the compensating vector field $V_{\mu}$ mentioned above, since whereas the isometry is broken by the background, the diffeomorphism invariance remains unbroken. Indeed, since only the scalar part of the metric fluctuation is relevant, the curvature perturbation is a separate object from the graviton as a transverse traceless tensor fluctuation which is still gauge invariant and massless. As the mass of the curvature perturbation is well below the inverse of the horizon radius, we can construct the effective field theory of inflation for it as well as the graviton to describe the cosmological perturbations in the inflationary era [38, 39].
@@ 3.1 Timelike isometry
$\begin{split}-\int dt&H_{1,M}=-\int d^{4}x\sqrt{\gamma}N{\cal H}_{M}\Big|_{\rm 1st}\\ &=\int d^{4}x\Big[\frac{\kappa a}{6}\Big(-\frac{3}{2}\frac{P_{\phi}^{2}}{a^{4}}-3a^{2}V\Big)h+\Big(-\frac{P_{\phi}}{a^{3}}\Big)\pi_{\phi}-\Big(a^{3}\frac{dV}{d\phi_{0}}\Big)\varphi+\kappa n\Big(-\frac{P_{\phi}^{2}}{2a^{3}}-a^{3}V\Big)\Big)\Big],\end{split}$ (23)
where indices are raised and lowered by $\delta^{ij}$ and $\delta_{ij}$, respectively (for example, $h$ indicates $\delta^{ij}h_{ij}$). We note that the last surface term in $\int dtH_{1,G}$ originates from the 3-dimensional Ricci tensor term $1/(2\kappa^{2})]R^{(3)}$ in (6). It is of ${\cal O}(\kappa^{-1})$ since we take a term linear in $\kappa h_{ij}$ of the Ricci tensor term containing a factor $\kappa^{-2}$. As commented in the Introduction, if the background geometry has the boundary, this surface term may not vanish, then for gauge invariance, the local operator which transforms nontrivially under the diffeomorphism can be dressed by the Wilson line, the nonlocal operator of ${\cal O}(\kappa)$ which does not vanish at the boundary [5] (see also discussion in [21]). However, (quasi-)dS space we are considering does not have a boundary, so the surface term vanishes. Then we cannot construct the gauge invariant operator by the nonlocal dressing. Instead, as the dS isometry is spontaneously broken, $\dot{\phi_{0}}$ (hence $\dot{H}$) does not vanish, and this leads to another way to construct the gauge invariant operator which is now local.
To see this, we notice from the expression for the linear order action that unlike the previous case where the timelike Killing vector exists and defines the time direction, each coefficient of the fluctuations ($\pi$, $h$, $\pi_{\phi}$, $\varphi$) except for $n$ in the linear order Hamiltonian $H_{1,G}+H_{1,M}$ does not vanish.  [^4] Of course, the total linear order Hamiltonian $H_{1,G}+H_{1,M}$ vanishes since it is a constraint, providing the relations between fluctuations. Since the total linear order action $S_{1,d}-\int dt(H_{1,G}+H_{1,M})$ is equivalent to the variation of the action around the classical solutions to the equations of motion, it must vanish, but this is achieved by the cancellation between coefficient in $S_{1,d}$ and that in the Hamiltonian part for each fluctuation. For instance, the coefficient of $a/(6\kappa)]h$ in $S_{1,d}$ given by $\dot{P}+HP+(3\kappa^{2}P_{\phi}/a)\dot{\phi_{0}}$ and the corresponding one in the Hamiltonian part given by $P^{2}/(12a^{2})]-\kappa^{2}[3P_{\phi}^{2}/(2a^{4})]-3\kappa^{2}a^{2}V$ are added to vanish, which is consistent with the equations of motion given by (17). In the same way, the coefficients of $\pi$ and $\pi_{\phi}$ in the Hamiltonian part are nothing more than the minus of the corresponding coefficients in $S_{1,d}$ :
@@
We can also interpret the curvature perturbation as the promotion of the scalar field $\phi(t,\mathbf{x})$ which is local but not gauge invariant to the local gauge invariant operator. That is, since the fluctuation $\zeta$ transforms as $\zeta\to\zeta+(H/\kappa)\epsilon^{0}\delta^{\mu}_{0}$ under the diffeomorphism $x^{\mu}\to x^{\mu}+\epsilon^{\mu}$, the scalar field becomes the gauge invariant operator when we replace $t$ by the gauge invariant combination $t-(\kappa/H)\zeta$ as (see also [27])
@@
Finally, we note that since the spontaneously broken dS isometry is a combination of the shift of $t$ and the rescaling of $r$, the fluctuation $\varphi$ can be interpreted as the Goldstone boson associated with the dilatation of the spatial coordinates, as well as that associated with the time translation. Moreover, another dS isometry, the special conformal invariance with three generators is also spontaneously broken [57], but unlike quantum field theory in Minkowski space the number of Goldstone bosons associated with the spontaneously broken symmetries is just one, rather than four. In fact, for a generic background, the number of Goldstone modes does not always coincide with the number of broken symmetries [58, 59].
@@ 3.2 Spacelike isometry
In fact, this confusion can be resolved by noticing that the transverse tracelss tensor fluctuation $h_{ij}^{TT}$ which is identified with the graviton is a gauge invariant operator. This is true in the local flat space (see footnote 5), and also can be found by comparing the covariant decomposition of the metric [60]
@@ 4 Accumulated fluctuation of the spacetime point
That is, as $d$ becomes larger than $4$, m=d2m=d-2 also increases as well. Therefore, so far as the expression (59) is not changed in higher dimensional spacetime except for the increase in the value of $n$, the fluctuation $\delta t/R$ decreases at late stage of the black hole evaporation, which allows the reliable construction of the island.
@@ 5 Conclusion
As we have investigated, when the isometry is spontaneously broken, the local gauge invariant operators can be constructed without breaking the diffeomorphism invariance. This can be achieved by the Stückelberg mechanism, where the fluctuation of the metric in the direction of the broken isometry transformation is combined with that of the matter field whose classical solution breaks the isometry to become physical. These operators also can be interpreted as the gauge invariant promotion of the local operators by introducing the ‘clock’ and the ‘rod’, which distinguish some particular time or spatial slice from another. Indeed, the concept of the clock and the rod has been employed to describe the thermodynamic properties of quantum gravity in terms of the von Neumann algebra [85, 86, 87, 88, 23] (for reviews on von Neumann algebra in quantum gravity, see, e.g., [89, 90]).
The construction of the local gauge invariant operator described above shows the tension between the locality and the isometry. Such a tension also can be found when we consider the effects of the accumulated fluctuation of the spacetime point generated by the frozen modes in the presence of the horizon. Since this fluctuation increases as $t^{1/2}$, the local region like the island may not be well constructed even if the local gauge invariant operators exist. In quasi-dS space describing the inflationary cosmology, the fluctuation of the spacetime point can be suppressed when the background is significantly deformed from dS space such that the timelike isometry is strongly broken. The structure of the local gauge invariant operator suggests that the similar situation can be found in the black hole interior after the Page time. In this case, the strong deformation of the background geometry may not be enough to suppress the fluctuation against its increasing behavior $\sim t^{1/2}$, unless the black hole can see the extra dimensions at late stage of the evaporation.

### v2

@@ 1 Introduction
Local field operators are the main building blocks of local quantum field theory, playing a key role in successfully reconciling quantum mechanics with special relativity (see, e.g., [2, 3] and references therein). Nevertheless, not all of them are physically meaningful when gravity is taken into account : for spacetime to be treated as a dynamical entity in a consistent way, physical operators must be invariant under the diffeomorphism.  [^1] See, e.g., [4, 5] for recent formal treatments. Indeed, it has been argued that such a gauge invariant operator is typically non-local, which is realized by ‘dressing’ the local operator at some point $x$ with a Wilson line connecting $x$ to the boundary [6, 7, 8, 9, 10, 11, 12]. This ‘dressing theorem’ can be shown by observing that if the background has a boundary or an asymptotic region, the diffeomorphism generator $\Phi$ contains a boundary term $\Phi_{\partial}$ of ${\cal O}(\kappa^{-1})$ where $\kappa=(8\pi G)^{1/2}$, thus can be written in the form of $\Phi=\Phi_{\rm bulk}+\Phi_{\partial}$. The local operator $O$ in the bulk commutes with the boundary term ($O,\Phi_{\partial}]=0$) while the commutator $O,\Phi_{\rm bulk}]$ of ${\cal O}(\kappa^{0})$ is typically nonzero. Then $O$ can be promoted to the gauge invariant operator by corrected by the Wilson line $\kappa O^{\prime}$ connecting the bulk and the boundary, as the commutator $\kappa O^{\prime},\Phi_{\partial}]$ of ${\cal O}(\kappa^{0})$ is nonvanishing but identical to $O,\Phi_{\rm bulk}]$. In the end, the dressed operator $O+\kappa O^{\prime}$ is gauge invariant and nonlocal.
@@
However, if all the gauge invariant operators are nonlocal, it becomes challenging to explain several quantum gravity phenomena. First, it was recently suggested that the black hole information paradox can be resolved by the appearance of a local region called the ‘island’ in the black hole interior after the Page time [13, 14, 15, 16, 17] (see also [18, 19] for reviews). As the operators localized on the island can be reconstructed from the operators for the Hawking radiation far away from the black hole, the entropy of the Hawking radiation does not increase indefinitely, which is consistent with unitarity. But in the absence of local gauge invariant operator, the operators localized on the island are difficult to be well defined. Second, whereas de Sitter (dS) space well describes the observed accelerating expansion of our universe [20, 21], it does not have a boundary, so the Wilson line connecting the bulk and the boundary is not defined.
@@
Regarding the first issue of the island, it has been suggested to construct the local operator by explicitly breaking the diffeomorphism invariance, resulting in a massive graviton [22, 23, 24, 25, 26]. This is motivated by the fact that when we consider the island of the anti-de Sitter (AdS) black hole in the context of AdS/CFT correspondence [27, 28, 29], the boundary of AdS space is assumed to be attached to the ‘bath’ in which gravity is not dynamical and the old Hawking radiation is contained. Whereas it is a convenient way to separate the Hawking radiation from gravitational dynamics, since energy is allowed to pass through the border between AdS space and the bath, the gravitational energy-momentum tensor is not conserved. In other words, the diffeomorphism invariance is broken by the boundary conditions. This can be modelled by introducing the compensating vector field $V_{\mu}$ which transforms as $V_{\mu}\to V_{\mu}-\epsilon_{\mu}$ under the diffeomorphism $x^{\mu}\to x^{\mu}+\epsilon^{\mu}$ before being fixed to some function of $x^{\mu}$. Then $V_{\mu}$ behaves like the Goldstone vector field : it combines with the metric fluctuation $h_{\mu\nu}$ to form the would-be gauge invariant combination $\tilde{h}_{\mu\nu}=h_{\mu\nu}+\nabla_{\mu}V_{\nu}+\nabla_{\nu}V_{\mu}$, from which the graviton mass term $(M^{4}/4)(\tilde{h}_{\mu\nu}\tilde{h}^{\mu\nu}-\tilde{h}^{\mu}_{\mu}\tilde{h}^{\nu}_{\nu})$ can be written in the Lagrangian density [26]. This corresponds to the imitation of (the tensor version of) the Stückelberg (or the Higgs in a broad sense) mechanism, in the vector version of which gauge invariance is spontaneously broken as the gauge boson becomes massive by absorbing a scalar transforming nonlinearly under the gauge transformation. Moreover, any local operator $O(x)$ can be promoted to the local would-be gauge invariant operator O(xV(x))O(x-V(x)) which is expected to be the form of operators localized on the island [30].
@@
Meanwhile, it was pointed out in [31] that the tension between the gauge invariance and the locality discussed so far arises only if the background has a symmetry called isometry. To see this intuitively we observe that the nonlocal connection between the bulk and the boundary is imposed by the isometry. This may be understood from the fact that any isometry breaking fluctuation raises the energy, which must be reflected in the fluctuation of the asymptotic charge like the Arnowitt-Deser-Misner (ADM) Hamiltonian. In contrast, in the absence of the isometry, the asymptotic charge may remain unaffected by the fluctuation in the bulk. Technically, the nonexistence of the local operator on the background having the isometry can be explained by the absence of the clock and the rod. To see this more concretely, suppose the background has a timelike isometry. In this case, one finds a time direction along which the metric does not change, hence the classical solution for the matter field is independent of the corresponding time coordinate $t$ as well. Then there is no way to distinguish some particular time slice from another, which is said to be the absence of the clock. In the same way, the spacelike isometry prevents distinguishing a particular spatial slice from another, then we say that the rod does not exist.
@@
In fact, the idea that the local gauge invariant operator can be constructed by combining the matter and the metric fluctuations goes back to [32] (see also discussion in [11, 12] which distinguishes the ‘gravitational dressing’ to which the dressing theorem is applied from ‘the field dressing’ where matter field also contributes to the dressing for gauge invariance). It was immediately used to describe the cosmological fluctuations around the Friedmann-Robertson-Walker (FRW) background which may not have a boundary, resolving the second issue [33, 34, 35, 36, 37, 38]. In particular, the curvature perturbation, the local gauge invariant operator made up of the scalar fluctuation and the trace of the spatial metric fluctuation, well describes the primordial inhomogeneity generated in the inflationary era, in which the background geometry is given by quasi-dS space where the dS isometry is slightly broken [39, 40]. The Stückelberg mechanism for the curvature perturbation must be distinguished from the imitated Stückelberg mechanism with the compensating vector field $V_{\mu}$ mentioned above, since whereas the isometry is broken by the background, the diffeomorphism invariance remains unbroken. Indeed, since only the scalar part of the metric fluctuation is relevant, the curvature perturbation is a separate object from the graviton, a transverse traceless tensor fluctuation of the metric which is gauge invariant. In this case, the graviton remains massless. As the mass of the curvature perturbation is well below the inverse of the horizon radius, we can construct the effective field theory of inflation for it as well as the graviton to describe the cosmological perturbations in the inflationary era [41, 42].
@@
We finally note that while [56] also claimed that the locality can be compatible with diffeomorphism invariance hence the island is well defined, the meaning of the locality there differs from ours. More concretely, the locality in [56] refers to the situation that the field configurations in two regions do not affect each other. This can be achieved even at the classical level when the perturbation restricted to the local region solves the linearized constraint equations without disturbing field values in other region. In contrast, the locality in our work is a quantum mechanical condition on operators, the commutativity of operators on different space-like separated region. We do not need to fix the field value in some region : the field can fluctuate in any spacetime point, which is nothing more than the quantum behavior of local operator. In this case, the locality seems to be implemented by describing all the gravitational phenomena including the island with the local operators, but it is not enough. We need to see whether the accumulated fluctuation of spacetime point can be sufficiently suppressed.
@@ 3.1 Timelike isometry
$\begin{split}-\int dt&H_{1,M}=-\int d^{4}x\sqrt{\gamma}N{\cal H}_{M}\Big|_{\rm 1st}\\ &=\int d^{4}x\Big[\frac{\kappa a}{6}\Big(-\frac{3}{2}\frac{P_{\phi}^{2}}{a^{4}}-3a^{2}V\Big)h+\Big(-\frac{P_{\phi}}{a^{3}}\Big)\pi_{\phi}-\Big(a^{3}\frac{dV}{d\phi_{0}}\Big)\varphi+\kappa n\Big(-\frac{P_{\phi}^{2}}{2a^{3}}-a^{3}V\Big)\Big],\end{split}$ (23)
where indices are raised and lowered by $\delta^{ij}$ and $\delta_{ij}$, respectively (for example, $h$ indicates $\delta^{ij}h_{ij}$). We note that the last surface term in $\int dtH_{1,G}$ originates from the 3-dimensional Ricci tensor term $1/(2\kappa^{2})]R^{(3)}$ in (6). It is of ${\cal O}(\kappa^{-1})$ since we take a term linear in $\kappa h_{ij}$ of the Ricci tensor term which contains a factor $\kappa^{-2}$. As commented in the Introduction, if the background geometry has the boundary, this surface term may not vanish, then for gauge invariance, the local operator which transforms nontrivially under the diffeomorphism can be dressed by the Wilson line, the nonlocal operator of ${\cal O}(\kappa)$ which does not vanish at the boundary [7] (see also discussion in [24]). However, (quasi-)dS space we are considering does not have a boundary, so the surface term vanishes. Then we cannot construct the gauge invariant operator by the nonlocal dressing. Instead, as the dS isometry is spontaneously broken, $\dot{\phi_{0}}$ (hence $\dot{H}$) does not vanish, and this leads to another way to construct the gauge invariant operator which is now local.
To see this, we infer from the expression for the linear order action that unlike the previous case where the timelike Killing vector exists and defines the time direction, each coefficient of the fluctuations ($\pi$, $h$, $\pi_{\phi}$, $\varphi$) except for $n$ in the linear order Hamiltonian $H_{1,G}+H_{1,M}$ does not vanish.  [^5] Of course, the total linear order Hamiltonian $H_{1,G}+H_{1,M}$ vanishes since it is a constraint, providing the relations between fluctuations. Since the total linear order action $S_{1,d}-\int dt(H_{1,G}+H_{1,M})$ is equivalent to the variation of the action around the classical solutions to the equations of motion, it must vanish, but this is achieved by the cancellation between coefficient in $S_{1,d}$ and that in the Hamiltonian part for each fluctuation. For instance, the coefficient of $a/(6\kappa)]h$ in $S_{1,d}$ given by $\dot{P}+HP+(3\kappa^{2}P_{\phi}/a)\dot{\phi_{0}}$ and the corresponding one in the Hamiltonian part given by $P^{2}/(12a^{2})]-\kappa^{2}[3P_{\phi}^{2}/(2a^{4})]-3\kappa^{2}a^{2}V$ are added to vanish, which is consistent with the equations of motion given by (17). In the same way, the coefficients of $\pi$ and $\pi_{\phi}$ in the Hamiltonian part are nothing more than the minus of the corresponding coefficients in $S_{1,d}$ :
@@
We can also interpret the curvature perturbation as the promotion of the scalar field $\phi(t,\mathbf{x})$ which is local but not gauge invariant to the local gauge invariant operator. That is, since the fluctuation $\zeta$ transforms as $\zeta\to\zeta+(H/\kappa)\epsilon^{0}\delta^{\mu}_{0}$ under the diffeomorphism $x^{\mu}\to x^{\mu}+\epsilon^{\mu}$, the scalar field can be promoted to the gauge invariant operator when we replace $t$ by the gauge invariant combination $t-(\kappa/H)\zeta$ as (see also [30])
@@
Finally, we note that since the spontaneously broken dS isometry is a combination of the shift of $t$ and the rescaling of $r$, the fluctuation $\varphi$ can be interpreted as the Goldstone boson associated with the dilatation of the spatial coordinates, as well as that associated with the time translation. Moreover, another dS isometry, the special conformal invariance with three generators is also spontaneously broken [61], but unlike quantum field theory in Minkowski space, the number of Goldstone bosons associated with the spontaneously broken symmetries is just one, rather than four. In fact, for a generic background, the number of Goldstone modes does not always coincide with the number of broken symmetries [62, 63].
@@ 3.2 Spacelike isometry
In fact, this confusion can be resolved by noticing that the transverse tracelss tensor fluctuation $h_{ij}^{TT}$ which is identified with the graviton is a gauge invariant operator. This is true in the local flat space (see footnote 6), and also can be found by comparing the covariant decomposition of the metric [64]
@@ 4 Accumulated fluctuation of the spacetime point
That is, as $d$ becomes larger than $4$, m=d2m=d-2 also increases as well. Therefore, so far as the expression (59) is not changed in higher dimensional spacetime except for the increase in the value of $n$, the fluctuation $\delta t/R$ decreases at late stage of the black hole evaporation, which allows the reliable construction of the local region at late state of the black hole evaporation.
But even in this case, there is another challenge for construction of the island described by local operators only, which comes from the fact that the island solution relies on the nonlocal nature of quantum gravity. More concretely, in order for the island to resolve the black hole information paradox, operators on the island must contain information on the Hawking radiation far away outside the black hole horizon. But if the operators on the island are local, it is not clear at present that how such nonlocal feature can be realized. Regarding this, the recent studies [89, 90] claim that when diffeomorphism (rather than isometry) is broken by the boundary condition (where the boundary refers to that between the bulk and the bath discussed in Introduction), the massive graviton encodes information on the heat bath. Then the compensating vector field plays the role of the wormhole operator connecting the island and the bath. Whether the similar mechanism can be realized when the isometry is spontaneously broken (without breaking diffeomorphism) is a subject of future investigation.
@@ 5 Conclusion
In this article, we have investigated the possibility of locality without breaking diffeomorphism invariance. Here by locality we mean the commutativity of operators on different space-like separated region. When the isometry is spontaneously broken, the local gauge invariant operators can be constructed without breaking the diffeomorphism invariance. This can be achieved by the Stückelberg mechanism, where the fluctuation of the metric in the direction of the broken isometry transformation is combined with that of the matter field whose classical solution breaks the isometry to become physical. These operators also can be interpreted as the gauge invariant promotion of the local operators by introducing the ‘clock’ and the ‘rod’, which distinguish some particular time or spatial slice from another. Indeed, the concept of the clock and the rod has been employed to describe the thermodynamic properties of quantum gravity in terms of the von Neumann algebra [91, 92, 93, 94, 26] (for reviews on von Neumann algebra in quantum gravity, see, e.g., [95, 96]).
The construction of the local gauge invariant operator described above shows the tension between the locality and the isometry. Such a tension also can be found when we consider the effects of the accumulated fluctuation of the spacetime point generated by the frozen modes in the presence of the horizon. Since this fluctuation increases as $t^{1/2}$, the local region like the island may not be well constructed even if the local gauge invariant operators exist. In quasi-dS space describing the inflationary cosmology, the fluctuation of the spacetime point can be suppressed when the background is significantly deformed from dS space such that the timelike isometry is strongly broken. The structure of the local gauge invariant operator suggests that the similar situation can be found in the black hole interior after the Page time. In this case, the strong deformation of the background geometry may not be enough to suppress the fluctuation against its increasing behavior $\sim t^{1/2}$, unless the black hole can see the extra dimensions at late stage of the evaporation. This condition on the accumulated fluctuation provides a stringent constraint for the appropriate implementation of the island, in addition to the existence of the local operator.